# Let’s Get to the Point: LLM-Supported Planning, Drafting, and Revising of Research-Paper Blog Posts

### v2

@@ 2.4. Human-LLM Scientific Writing
Several works have investigated how to support human-LLM writing. Most of these works focus on tools for creative writing (Singh et al., 2022; Chung et al., 2022; Mirowski et al., 2023), but a few cater to scientific writing. A few works have investigated human-LLM scientific writing for broader science communication tasks that do not involve a source document. Gero et al. present a system for generating scientific “sparks”, or inspiring sentences (Gero et al., 2022). Long et al. examine how LLM scaffolding can help people generate relatable hooks for complex scientific topics (Long et al., 2023), and Kim et al. investigate how science writers can generate extended metaphors for scientific ideas with the help of an LLM (Kim et al., 2023). The few works that have explored human-AI scientific summarization do not allow for in-context selection of source content but rather allow users to select concepts (Zhang et al., 2023b), out-of-context sentences extracted based on queries (Bayatmakou et al., 2022), suggested summary content (Lim et al., 2024). Lim et al.’s Q-and-A research-paper summarization tool AQUA also allows users to prompt the LLM with relevant questions (Lim et al., 2024). Meanwhile, Google’s tool NotebookLM for general human-AI document summarization does not permit any content selection (Google, [n. d.]).
Based on prior work (Mehlenbacher, 2019), popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), and feedback from three communications experts, we determined a set of guidelines for writing research-paper blog posts, which indicate that the shortest blog posts are recommended to be around 500 words in length, making blog posts a long-form output. Because research-paper blog posts provide long-form scientific communication, drafting them requires LLM instructions unique to common blog post sections, and revising them requires balancing the writing style for accuracy, engagement, ease of understanding, and concision. Most of the previously cited forms of human-LLM scientific writing above are the length of a sentence or paragraph. The exceptions are AQUA’s Q-and-A paper summary (Lim et al., 2024) and NotebookLM’s many summary forms (e.g., FAQ, study guide, briefing doc, podcast). However, AQUA’s scaffolding is limited to suggested questions and answers as well as a space for user-provided additional prompting, and NotebookLM provides no scaffolding to support customization of their document summaries. Furthermore, our work is unique in its dedication to supporting mixed-initiative writing of research-paper blog posts.
@@ 2.5. Scientific Blog Posts
Scientists can engage with broad community through various channels, including press releases, magazines, journals, and tweets (August et al., 2020; Gero et al., 2022). This work focuses on scientific blog posts, which offer increased recognition and transparency of one’s work, writing practice, editorial freedom, and public engagement (Jarreau, 2015; Blanchard, 2011; Zou and Hyland, 2019). There are several forms of scientific blog posts, such as academic commentary, free commentary, and mediation of research to laypersons (Mahrt and Puschmann, 2014). Here, we focus on blog posts with the primary goal of communicating information from a research paper to people unfamiliar with the paper’s specific research topic.
Figure 2. Papers-to-Posts’ user interface (1/2) showing the Plan and Draft functionality. a) The Blog-Post Area, where the user writes the blog post in sections. b) The Planning-Support Area, which contains both the interactive paper outline and original paper text. c) The Drafting-Support Area, which contains inputs and outputs for generating text. (Not Pictured: The generation output appears below the Regenerate button in the bottom right, with a button to copy the output text and buttons to view previous texts and their inputs.)
@@ 3. The Papers-to-Posts System
We developed Papers-to-Posts based on key design goals and pilot study feedback, which helped us to identify system features to add for increased transparency, flexibility, and ease of use. This section describes the design goals for Papers-to-Posts and provides a walkthrough of the resulting system’s four main areas. [^1] A walkthrough video of the tool may be found in the supplementary materials.
@@ 3.1. Design Goals
We determined the design goals for Papers-to-Posts based on guidelines for academic blog posts from popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), prior work (Mehlenbacher, 2019), and feedback from three communications experts, recruited from our institutions. There are multiple guidelines on the content and format for different sections of the academic blog post (e.g., introduction, results) (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Mehlenbacher, 2019). With respect to the blog post’s tone, guidelines suggest avoiding jargon (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), sensationalism (noted by a communications expert), and wordiness (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]).
The resulting design goals are:
  • DG1: Help researchers to plan what paper content to include in their blog post.
  • DG2: Help researchers to draft their content in sections amenable to a research-paper blog post.
  • DG3: Help researchers to revise the blog post’s tone to be scientifically accurate but also concise, easy to understand, and engaging.
3.2. Walkthrough of Start-Up Step
3.2.1. Frontend
In the Blog Post Area (Figure 2a), the user starts by reviewing the initial blog post draft. The blog post draft is divided into sections. The user clicks on whichever section they want to work on first, and the tool highlights that section for reference. At the top, the user views the blog post’s word count, a preview of the full blog post, and tips for writing blog posts, which are based on popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), prior work (Mehlenbacher, 2019), and feedback from three communications experts.
3.2.2. Backend
In order to obtain the initial blog post draft for a given research paper, we completed a few steps (Figure 1a). In the first step, we used a public tool for converting scientific PDFs to HTML [^2] https://papertohtml.org to extract the paper text from the paper PDF. For the lab study, we then manually adjusted the parsed HTML so that the section headers were correct, no section or subsection was completely missing, and only the main paper text was included (i.e., no appendix, references, or footnotes). We also manually retrieved the tables and figures from the parsing. For the real-world deployment study, the parsing was automatic to better align with a real-world setting. For the second step, we used GPT-3.5 Turbo to generate between one and three bullet points to summarize each paragraph in the parsed paper, depending on the length of the paragraph. While we used GPT-4 (GPT-4-0125-preview in the deployment study) for all of Papers-to-Posts’s other tasks, we decided to use GPT-3.5-Turbo for this task for two reasons: 1) it produced shorter bullet points that were quicker to read, and 2) it was many times cheaper for this token-intensive generation task. All together, the generated bullet points formed a paper outline. For the third step, we provided the LLM with all of the paper’s bullet points along with the request to select 10 bullet points relevant to the initial four blog post sections: introduction, methods, results, and conclusion. We requested 10 bullet points per blog post section in order to provide the model sufficient content without overwhelming it. For the fourth step, we prompted the LLM to generate each section in succession using the selected bullet points for that section, the paragraphs associated with the selected bullet points (to better ground the model), the portion of the blog post that had already been generated, and section-specific guidelines based on popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), prior work (Mehlenbacher, 2019), and feedback from three communications experts. We observed that this process of generating one draft section at a time with the context of prior sections helped to make the draft more coherent while avoiding overwhelming the model with instructions for multiple sections. The model was prompted to make each section around 125 to 250 words and around one to three paragraphs. The prompts for generating the initial draft may be found in Appendix B.2. An example initial draft and associated selected bullet points is in Appendix F.1 and F.2.
@@ 3.3. Walkthrough of Planning Step
3.3.1. Frontend
In the Planning-Support Area (Figure 2b), the user works with the tool to plan the blog post section on which they are currently working (DG1). This area has two tabs labeled Paper Outline and Paper Text. Under the Paper Outline tab, the user views the paper’s outline, divided into sets of one to three bullet points; each set is associated with a paragraph from the original text. The user views a bullet point set’s corresponding paragraph by clicking the button to the left, and the reverse is true under the Paper Text tab, where the user views the original paper text divided into paragraphs. Under both tabs, the information is divided into collapsible sections based on the paper’s highest-level sections. [^3] In the lab study, the tool presented the figures and tables in a final collapsible section. Their captions could be copied for pasting in the editor to help the user plan where to put figures in the final blog post. We did not include this feature in the deployment study, as the automatic paper parsing did not extract this information.
Under the Paper Outline tab, the user reviews the tool’s pre-selected bullet points for the current blog post section by scrolling through the full outline or clicking the “Show Selected Only” button below the paper area. When this filter button is activated and the user is under the Paper Text tab, if a paragraph is not selected but a corresponding bullet point is selected, the tool notes that “A corresponding bulletpoint(s) is still selected”; the opposite is true for the Paper Outline tab. The user clicks the “Close All Paper Sections” button below the paper area if they want to get a higher-level view of the paper. The user then adjusts the selection of bullet points or paragraphs for the LLM to use in generating text for the current blog post section. Selecting a paragraph does not affect the selection of associated bullet points, and vice versa. If the user wants to reset the content selection, they click the “Clear All Selected” button below the paper area.
3.3.2. Backend
As shown in Figure 1b, all planning actions (as well as drafting and revising actions) are connected to the user’s selected blog post section. In other words, the state and history of all inputs and outputs for text generation are stored separately for each section. This is somewhat similar to the fragmented document history proposed by Buschek (Buschek, 2024).
@@ 3.4. Walkthrough of Drafting Step
3.4.1. Frontend
@@
In the Drafting-Support Area (Figure 2c), the user works with the tool to synthesize text for the blog post section on which they are currently working (DG2). The user reviews and adjusts the LLM’s section-specific instructions for generating text. At the top of the Drafting Support Area, the user changes the desired length of the generated text to their liking. The length defaults to “auto” and may be updated to one sentence, one paragraph, or a few paragraphs. Below, the user notes the toggle that is turned on by default. This toggle lets the user know that the context of blog post sections above the highlighted one is provided to the LLM in order to generate text for this section, and they may turn this toggle off if they do not want this context to influence the generated text for this section. In the “Custom Bulletpoints” text field, the user enters any additional bullet points that they want their generated text for the highlighted section to cover. In the “Custom Instructions” text field, the user adjusts the instructions for the highlighted section’s generated text. For each of the initial blog post sections, this box is pre-filled with section-specific instructions, which are based on the blog post guidelines we collected. In the text field labelled “Starting Text to Continue from”, the user optionally provides text that should be at the start of the generated text. Finally, the user clicks the “generate” button to generate text for the highlighted section, which appears below. The user then copies and pastes the generated text into the highlighted blog post section and edits it as they see fit. The user may peruse a history of their content selection, LLM instructions, and text generations for the highlighted section.
The user also drafts text for the blog post by adding a new section. In the Blog Post Area, the user clicks the plus-sign button next to an existing blog post section, and the tool places the new section below the existing one. The tool gives the user the option to either create a blank section or generate a section based on the section header that they provide. If the user decides the new section should be in another location within the blog post, they use the section’s associated up and down arrow buttons to move the section to the desired point. For this new section, the user may return to the planning step or adjust the LLM’s instructions for generating text. If the user decides not to include the section in the blog post, they delete it using the trash button.
3.4.2. Backend
@@
As shown in Figure 1c, the LLM (GPT-4 in lab study; GPT-4-0125-preview in deployment) generates text for the highlighted blog post section with the user’s adjusted selection of bullet points and paragraphs as well as their adjusted section-specific instructions (described in Section 3.4.1). With respect to the selected bullet points, the model receives the context of the associated paragraph for each bullet point selected, even if the associated paragraph was not selected by the user. This is to ensure the model’s generation is grounded in the actual paper text. When the user desires the “auto” length for generated text, the model is instructed to generate text that is 125 to 250 words and between one and three paragraphs. To generate a new section, we prompt the LLM to select 10 bullet points for the section based on the provided header and to generate text focused on those bullet points. Again, the model is provided the context of the bullet points’ associated paragraphs. The prompts for drafting text may be found in Appendix B.4, and example outputs are in Appendix F.1 and F.2.
Figure 3. Papers-to-Posts’s user interface (2/2), showing the Revise functionality. d) The Revising-Support Area, which contains inputs and outputs for modifying text and is located below the Drafting-Support Area. (Not Pictured: The modification output appears below the modification buttons, with a button to copy the output text and buttons to view previous texts and their inputs.)
[…]

### v3

@@ 2.3. Human-LLM Scientific Writing
Several works have investigated how to support human-LLM writing. Most of these works focus on tools for creative writing (Singh et al., 2022; Chung et al., 2022; Mirowski et al., 2023), but a few cater to scientific writing. A few works have investigated human-LLM scientific writing for broader science communication tasks that do not involve a source document. Gero et al. present a system for generating scientific “sparks,” or inspiring sentences (Gero et al., 2022). Long et al. examine how LLM scaffolding can help people generate relatable hooks for complex scientific topics (Long et al., 2023), and Kim et al. investigate how science writers can generate extended metaphors for scientific ideas with the help of an LLM (Kim et al., 2023). The few prior works that have explored human-AI scientific summarization (Lim et al., 2024; Zhang et al., 2023b; Bayatmakou et al., 2022) do not support fully grounded and detailed summaries like Papers-to-Posts, as described in Section 2.1. Also of note, Google’s tool NotebookLM for general human-AI document summarization does not permit any content selection (Google, [n. d.]).
@@ 2.4. Scientific Blog Posts
Scientists can engage with broader communities through various channels, including press releases, magazines, journals, and tweets (August et al., 2020; Gero et al., 2022). This work focuses on scientific blog posts, which offer increased recognition and transparency of one’s work, editorial freedom, and public engagement (Jarreau, 2015; Blanchard, 2011; Zou and Hyland, 2019). There are several forms of scientific blog posts, such as academic commentary, free commentary, and mediation of research to laypersons (Mahrt and Puschmann, 2014). Here, we focus on blog posts with the primary goal of communicating information from a research paper to people unfamiliar with the paper’s specific research topic.
Figure 3. Papers-to-Posts’ user interface. a) The Blog-Post Area, where the user writes the blog post in sections. b) The Planning-Support Area, which contains both the interactive paper outline and original paper text. c1) The Drafting-Support Area, which contains inputs and outputs for generating text. c2) The Drafting-Support Area continued, seen if one scrolls below the c1 area. d) The Revising-Support Area, which contains inputs and outputs for modifying text and is located below the Drafting-Support Area. (Not Pictured: The modification output appears below the modification buttons, with a button to copy the output text and buttons to view previous modifications and their inputs.)
@@
There are two images next to each other. The left image has two panels. The left panel is labeled ”a” and titled ”Blog Post Draft.” It shows the blog post draft with the introduction section open and highlighted in light yellow, while the rest of the blog post sections are collapsed and have a white background. The draft shows a box for the title at the top followed by each initial draft section (introduction, methods, results, conclusion, call to action). Each section has a few buttons next to its header: a short arrow for collapsing/expanding the section, up and down arrows for moving the sections to a new location in the blog post, a trash icon, and a plus icon for adding a new section. Meanwhile, the right panel of the left image is light yellow and titled ”Generate writing for highlighted blog post section.” The top half of the panel is labeled ”b” and the bottom half is labeled ”c1.” The top half contains the instruction ”1. Select any paper paragraphs or bullet points to cover in generated writing.” Below, it shows two tab buttons labeled ”Paper Text” and ”Paper Outline.” The Paper Outline tab is clicked, and beneath there is a light gray box with paper section headers and a few checked bullet points under some of the headers. The last section header visible is cut off, indicating there are unseen sections. Underneath, there are three buttons: ”Clear All Selected,” ”Show Selected Only,” and ”Close All Paper Sections.” In the bottom half of the panel, from top to bottom, there is the following: 1) text saying ”2. Selected desired length of generated writing:” followed by a drop-down menu set to ”auto” to its right; 2) text saying ”3. Check if want generated writing to consider context of blog post sections above this one: followed by a checked box to its right”; 3) text saying ”4. Optionally enter any custom bullet points to cover in generated writing.” followed by an empty text box; 4) text saying ”5. Optionally enter any custom instructions for generated writing to follow.” followed by a text box with text saying ”-Present hook (e.g., interesting fact, quote, promise of change in knowledge, illustrating example of the topic). If context allows, visual or sensory”; 5) text saying ”6. Optionally enter any starting text from which the generated writing should continue.” The second image of the figure shows the continuation of the right panel (if one was to scroll down). The top half of the panel shown is labeled ”c2” and the bottom half is labeled ”d.” In the top half, from top to bottom, there is the following: 1) text saying ”6. Optionally enter any starting text from which the generated writing should continue.” followed by an empty text box; 2) a button labeled ”Generate”; 3) A paragraph of text with a light gray background that starts with ”In the rapidly evolving field of machine-learning, attention-based architectures…” and continues; 4) a button that says ”Copy Generation” followed by a left arrow, ”1/1” and a right arrow to its right. The bottom half of the panel is enclosed in a black border. It is titled ”Modify Writing.” From top to bottom, it contains the following: 1) text saying ”1. Enter writing to modify.*” followed by an empty text box; 2) text saying ”2. Select desired length of modified writing:” followed by a drop-down menu set to ”auto”; 3) text saying ”3. Check if want modified writing to consider context of blog post sections above this one:” followed by a checked box to its right; 4) text saying ”4. check if want modified writing to be based on selected paragraphs and bullet point:” followed by a checked box to the right; 5) text saying ”5. optionally enter any custom instructions for modified writing to follow.” followed by an empty text box; 6) text saying ”6. click modification to perform. (Custom Modification will be based on custom instructions above.); 7) Buttons labeled ”Expand,” ”Condense,” and ”Simpler Terms”; 8) Buttons labeled ”Less Dramatic,” ”More Dramatic,” and ”Custom Modification.”
@@ 3. The Papers-to-Posts System
This section describes the design goals for Papers-to-Posts and provides a walkthrough of the resulting system’s four main panel areas, which show the blog post draft and enable executing the Plan-Draft-Revise workflow on individual blog post sections. [^1] A walkthrough video of the tool may be found in the supplementary materials. We developed Papers-to-Posts based on key design goals and pilot study feedback. The pilot study feedback helped us to identify system features to add for increased transparency, flexibility, and ease of use.
@@ 3.1. Design Goals
When communicating scientific research in a detailed summary, making sure all important details are covered is key. Otherwise, the research may be misinterpreted, leading to misinformed readers. Indeed, in a closely related work on human-LLM scientific text summarization, a featured design goal based on their preliminary survey of professors was to provide provenance for summary content by showing “direct or indirect contributors to a summary to help the user verify whether the summary reflects the structure and key components of the original document” (Zhang et al., 2023b). This also applies to similar domains, such as law and business, in which delivering information with accuracy and precision is critical. However, existing text summarization tools (including the one just referenced) either 1) do not fully ground the summary content selection in the source, which makes understanding what content is versus is not included difficult, or 2) only support selection of summary content at the level of source sentences, which is unwieldy when writing a detailed summary (see Section 2.1).
@@
Furthermore, we identified and synthesized guidelines for academic blog posts from popular resources (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]), prior work (Mehlenbacher, 2019), and feedback from three communications experts, recruited from our institutions. There are multiple guidelines on the content, format, and language for academic blog posts. Papers-to-Posts seeks to address these guidelines by default but does not enforce them. [^2] The synthesized guidelines may be found in the appendix.
Thus, Papers-to-Posts’ design goals are as follows (science domain-specific details in parentheses):
  • DG1: Help users (researchers) to understand what long-document (research-paper) content is versus is not included in its detailed summary (blog post).
  • DG2: Help users (researchers) to control and iterate on what long-document (research-paper) content is included in a detailed summary (blog post).
  • DG3: Help users (researchers) to follow domain-specific guidelines (guidelines for academic blog posts), as a starting point.
3.2. Implementation Details
The frontend of Papers-to-Posts was developed using React and TypeScript, while the backend was developed with Python. Unless otherwise noted, the LLM was GPT-4 in the lab study and GPT-4-0125-preview in the deployment study.
3.3. Walkthrough of Warm Start Step
The Blog Post Area (Figure 3a) is the user’s main working area for reviewing the target blog post artifact. At the top, the user views the blog post’s word count, a preview of the full blog post, and tips for writing blog posts (described in Section 3.1). The user then reviews the blog post draft, divided into sections. To mitigate cognitive overload, the user selects one section at a time to focus on. The selected section is the target of subsequent Plan-Draft-Revise workflow actions, described in the other step descriptions below.
In order to obtain the initial blog post draft for the user’s input research paper, we developed an LLM-powered pipeline that first generates a reverse source outline, selects relevant points, and then generates a draft (Figure 1a).
3.3.1. Preprocessing
We used a public tool for converting scientific PDFs to HTML [^3] https://papertohtml.org to extract the paper text from the paper PDF. For the lab study, we then manually adjusted the parsed HTML so that the section headers were correct, no section or subsection was completely missing, and only the main paper text was included (i.e., no appendix, references, or footnotes). We also manually retrieved the tables and figures from the parsing. For the deployment study, the text parsing was automatic to better align with a real-world setting, and there was no table or figure retrieval.
3.3.2. Outline generation
We used GPT-3.5 Turbo to generate between one and three bullet points to summarize each paragraph in the parsed paper, depending on the length of the paragraph. While we used GPT-4 (GPT-4-0125-preview in the deployment study) for all of Papers-to-Posts’ other tasks, we decided to use GPT-3.5-Turbo for this task for two reasons: 1) it produced shorter bullet points that were quicker to read, and 2) it was many times cheaper for this token-intensive generation task. All together, the generated bullet points formed a paper outline.
3.3.3. Outline point selection
We provided the LLM with all of the paper’s bullet points along with the request to select 10 bullet points relevant to the initial four blog post sections: introduction, methods, results, and conclusion. We requested 10 bullet points per blog post section in order to provide the model sufficient content without overwhelming it.
3.3.4. Summary generation
We prompted the LLM to generate each section in succession using the selected bullet points for that section, the paragraphs associated with the selected bullet points (to better ground the model), the portion of the blog post that had already been generated, and section-specific guidelines (see Section 3.1). We observed that this process of generating one draft section at a time with the context of prior sections helped to make the draft more coherent while avoiding overwhelming the model with instructions for multiple sections. The model was prompted to make each section around 125 to 250 words and around one to three paragraphs. [^4] The prompts for generating the initial draft are in Appendix B.2. An example initial draft and associated selected bullet points are in Appendices F.1 and F.2.
@@ 3.4. Walkthrough of Planning Step
In the Planning-Support Area (Figure 3b), the user works with Papers-to-Posts’s interative reverse source outline to plan the blog post section on which they are currently working, addressing DG1 and DG2. All of the user’s planning actions (as well as drafting and revising actions) are connected to the user’s selected blog post section. In other words, the state and history of all inputs and outputs for text generation are stored separately for each section. This is somewhat similar to the fragmented document history proposed by Buschek (Buschek, 2024).
The user sees the interactive reverse source outline, divided into sets of bullet points, each corresponding to a paragraph from the original text, which can be viewed by clicking the button to the left of the bullet point set.  [^5] In the lab study, the tool presented figures/tables in a final collapsible portion of the paper outline. Their captions could be copied for pasting in the editor to help the user plan where to put figures in the final blog post. They review the tool’s pre-selected bullet points for the current blog post section by scrolling through the full outline or clicking the “Show Selected Only” button. [^6] When this filter button is activated and the user is under the Paper Text tab, if a paragraph is not selected but a corresponding bullet point is selected, the tool notes that “A corresponding bulletpoint(s) is still selected”; the opposite is true for the Paper Outline tab. The user adjusts the selection of bullet points (or paragraphs) for the LLM to use in generating text for the current blog post section (Figure 1b). Selecting a paragraph does not affect the selection of associated bullet points, and vice versa, from the user’s perspective. The system, however, does receive associated paragraphs with selected bullet points for grounding.
@@ 3.5. Walkthrough of Drafting Step
In the Drafting-Support Area (Figure 3c), the user works with the tool to synthesize text for the blog post section on which they are currently working, addressing DG2 and DG3.
3.5.1. Section Revision
@@
At the top of the Drafting Support Area, the user changes the desired length of the generated text to their liking. The length defaults to “auto” and may be updated to one sentence, one paragraph, or a few paragraphs. Below, the user sees the toggle is on to provide the LLM with the context of blog post sections above the one on which they are working. In the “Custom Bulletpoints” text field, the user enters any additional bullet points that they want their generated text for the blog post section to cover. In the “Custom Instructions” text field, the user adjusts the instructions for the section’s generated text. For each of the initial draft sections, this box is pre-filled with section-specific instructions, which are based on blog post guidelines we collected on content and formatting for different academic blog post sections (Kudos, [n. d.]; UWaterloo, [n. d.]; Patrick Dunleavy, [n. d.]; CMU, [n. d.]; Mehlenbacher, 2019). In the text field labeled “Starting Text to Continue from”, the user optionally provides text that should be at the start of the generated text.
@@
The LLM then generates revised text for the section at hand with the user’s adjusted selection of bullet points and paragraphs as well as their adjusted section-specific instructions (Figure 1b). [^7] The prompts for drafting text may be found in Appendix B.4, and example outputs are in Appendices F.1 and F.2. With respect to the selected bullet points, the model receives the context of the associated paragraph for each selected bullet point, even if the associated paragraph was not selected by the user. This is to ensure that the model’s generation is grounded in the actual paper text. When the user selects the “auto” length for generated text, the model is instructed to generate text that is 125 to 250 words and between one and three paragraphs.
The user copies and pastes all or part of the generated text into the highlighted blog post section and edits it as they see fit. If the user wants to remember what they generated previously for this section, they may peruse a history of their inputs (paper content selection, LLM instructions) and outputs (text generations).
3.5.2. Section Addition
The user can also add new sections to the blog post (Figure 1c). In the Blog Post Area, the user clicks the plus-sign button next to an existing blog post section, and the tool places the new section below the existing one. The tool gives the user the option to either create a blank section or generate a section based on the section header that they provide. To generate a new section, the LLM selects 10 bullet points for the section based on the provided header and generates text based on those bullet points and associated paragraphs.